Problem 8-29 Completing a Master Budget (LO8-2, LO8-4, LO8-7, LO8-8, LO8-9, LO8-10] The following data relate to the operatioComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 PreComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 Pre
Answer #1
1) Shilow company
Schedule of Expected cash collections
April May June Quarter
Cash sales 46200 49200 64200 159600
credit sales 24,400 30800 32800 88,000
total collections 70600 80000 97000 247600
Accounts receivable = 107000*40%= 42800
2) Merchandise purchase budget
April May June Quarter
Budgeted cost of goods sold 57750 61500 80250 199500
Add Desired ending inventory 49200 64200 34,800 34,800
total needs 106950 125700 115050 234300
less beginning inventory 46,200 49,200 64,200 46,200
Required purchases 60,750 76,500 50,850 188,100
3) Schedule of Cash disbursements-Merchandise purhcase
April May June Quarter
March purchases 27,675 27,675
April purchases 30375 30,375 60750
May purchases 38250 38,250 76500
June purchases 25425 25425
total disbursements 58,050 68625 63675 190,350
4) Cash budget
April May June Quarter
Beginning cash balance 8,600 4,290 4,505 8,600
Add Cash collectiosn 70600 80000 97000 247600
total cas h available 79,200 84,290 101,505 256,200
less cash disbursements
for inventory 58,050 68625 63675 190,350
for expenses 17260 18160 22660 58080
for equipment 2,600 0 0 2,600
total cash disbursements 77,910 86785 86335 251,030
Excess(Deficiency)of cash 1,290 -2,495 15,170 5,170
Financing:
Borrowings 3,000 7,000 0 10,000
Repayments 0 -10,000 -10,000
interest 0 -230 -230
total financing 1,000 7,000 -10230 -230
Ending cash balance 4,290 4,505 4,940 4,940
interest = 3000*1%*3= 90
7000*1%*2= 140
230
5) income statement
Sales 266000
cost of goods sold
Beginning inventor 46,200
Add purchases 188,100
goods available for sale 234,300
ending inventory 34,800 199,500
Gross margin 66,500
Selling and administrative expense
commissions 31920
rent (3400*3) 10200
Depreciation (891*3) 2673
other expenses 15960 60753
net operating 5,747
interest expense -230
net income 5,517
Balance sheet
Assets
current assets
Cash 4,940
Accounts receivable 42,800
inventory 34,800
total current assets 82,540
Building And equipment ,net 118727
total Assets 201,267
liabilities And stockholder 's Equity
Accounts payable 25,425
total current assets 25,425
Stockholder's Equity
Capital stock 150,000
Retained earnings 25,842 175,842
total liabilites & stockholders Equity 201,267